OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8>

OpenStudy (sirm3d):

the angle \(\theta\) between two vectors is \[\cos \theta=\frac{u\cdot v}{|u| |v|}\]

OpenStudy (anonymous):

okay thank you i just found it in my book